### v1

# An End-to-End Room Geometry Constrained Depth Estimation Framework for Indoor Panorama Images

Abstract
Predicting spherical pixel depth from monocular $360^{\circ}$ indoor panoramas is critical for many vision applications. However, existing methods focus on pixel-level accuracy, causing oversmoothed room corners and noise sensitivity. In this paper, we propose a depth estimation framework based on room geometry constraints, which extracts room geometry information through layout prediction and integrates those information into the depth estimation process through background segmentation mechanism. At the model level, our framework comprises a shared feature encoder followed by task-specific decoders for layout estimation, depth estimation, and background segmentation. The shared encoder extracts multi-scale features, which are subsequently processed by individual decoders to generate initial predictions: a depth map, a room layout map, and a background segmentation map. Furthermore, our framework incorporates two strategies: a room geometry-based background depth resolving strategy and a background-segmentation-guided fusion mechanism. The proposed room-geometry-based background depth resolving strategy leverages the room layout and the depth decoder’s output to generate the corresponding background depth map. Then, a background-segmentation-guided fusion strategy derives fusion weights for the background and coarse depth maps from the segmentation decoder’s predictions. Extensive experimental results on the Stanford2D3D, Matterport3D and Structured3D datasets show that our proposed methods can achieve significantly superior performance than current open-source methods. Our code is available at https://github.com/emiyaning/RGCNet.
I Introduction
With the advent of consumer-grade omnidirectional cameras such as the Ricoh Theta, Samsung Gear 360, and Insta360, the acquisition of panoramic images has been significantly simplified. The $180^{\circ}\times 360^{\circ}$ field of view offered by panoramas renders them particularly valuable for indoor 3D perception applications[1]. To enable 3D understanding of indoor environments from a single omnidirectional image, depth estimation becomes a fundamental requirement[2, 3, 4, 5, 6, 7, 8]. This potential for inferring entire scene structure from a single panorama has motivated active research on panoramic depth estimation.
Figure 1: 3D visualization of panorama depth estimator’s predictions on the Stanford2d3d dataset. The left side shows the ground-truth visualization, middle column shows the visualization of Panoformer’s prediction, right side shows the visualization of our framework’s prediction.
However, the ultra-wide field of view in panoramic imaging inherently introduces object distortion. To address this issue, current panoramic depth estimation approaches primarily employ two kinds of pathway[9]. One strategy designs dedicated feature extractors motivated by panoramic imaging principles to handle distortion directly[10, 11]. The other strategy projects the equirectangular panorama (ERP) into six cubemap faces, employs feature encoders separately on both the ERP and cubemap projections [12], fuses their features, and subsequently predicts depth. These existing methods have achieved remarkable achievements in the field of depth estimation of panoramic images. However, current panoramic depth estimation methods excessively prioritize local pixel accuracy over geometric room structures, leading to two issues: (1) Noisy ground truth data induces local overfitting that propagates errors to adjacent regions; (2) Local correlations cause inaccurate predictions at 3D discontinuity regions (e.g., wall corners). As described in Fig. 1, the pre-trained panorama depth estimator PanoFormer[10] shows limited performance on the wall corners while easily be disturbed by noisy ground truth data. These two problems caused by ignoring room geometry seriously hinder the application of existing depth estimation models to real-world indoor scenarios.
To address existing problems in panoramic depth estimation, this paper proposes a depth estimation framework named as room geometry constrained depth estimation network (RGCNet). The proposed RGCNet based on multi-task learning unifying three complementary tasks: depth estimation, room layout estimation, and background segmentation. This architecture employs a shared panoramic encoder to extract multi-scale features, with three dedicated decoders generating: coarse depth map, room layout map, and background segmentation mask from these features. The depth estimation branch retains the architecture of existing high-performance depth estimators [10, 11]. However, our RGCNet treats the output of the depth decoder as a first-stage prediction and refines it using the room geometry constrain which generated from room layout and background segmentation. In our framework, the initial depth estimate and room layout are jointly utilized to compute the background depth maps. By incorporating geometric constraints of the room, these derived background depth maps are not only more robust to noise in the ground-truth depth but also yield higher accuracy in structured regions such as walls and corners. Furthermore, RGCNet leverages an end-to-end multi-task framework to decode the background depth and fuse it with the coarse depth prediction in a single efficient step. To generate the background map, a proposed room-geometry-based background depth resolving strategy leverages the layout and depth decoder’s output to calculating the corresponding background depth. To fuse the background and depth decoder’s prediction, a background-segmentation-guided fusion strategy derives fusion weights for the background and coarse depth maps from the segmentation decoder’s predictions. In general, our contributions can be summarized into the following three points:
  • A room-layout constrained depth estimation framework RGCNet has been proposed to use room structural geometry to get corrected depth predictions.
  • A background depth map resolved strategy has been proposed, which extract the room structural geometry information from multi-task decoder’s prediction.
  • A adaptively fusion strategy has been proposed, which adaptively based on extracted the geometry information to refine the depth prediction.
II Related Works
A core challenge in panoramic depth estimation is image distortion. Existing research primarily employs three approaches: 1) estimating depth solely on a single projection; 2) projecting the panorama onto multiple modalities for depth estimation; and 3) using generated background depth to guide estimation.
II-A Single Projection Inputs
Panoramic images employ spherical representations with $180^{\circ}$ vertical and $360^{\circ}$ horizontal fields of view. Processing typically involves projecting these spherical images onto 2D planes through perspective mapping. Common modalities include equirectangular[13, 14, 10, 11, 15, 16, 17], tangent[18, 19, 20], and icosahedron projections[21, 22].
For depth estimation, equirectangular projection predominates. Omnidepth[13] introduced RectNet for efficient equirectangular feature extraction and depth prediction. ODE-CNN[14] proposed a hardware-software co-design system comprising: 1) a panoramic camera with binocular depth sensors, and 2) a depth estimation model leveraging RectNet baseline enhanced with spherical feature transformers between encoder-decoder stages and deformable convolutional spatial propagation for final prediction. Following vision transformers’ emergence, specialized models have proliferated: Panoformer[10] developed reference-point window self-attention for equirectangular features within an encoder-decoder architecture; Egformer[11], SGFormer[15], GLPanoDepth[23] subsequently incorporated global receptive fields and spherical geometry constraints. To address panoramic depth data scarcity, self-supervised methods [16, 17] operate directly on equirectangular RGB images.
Alternatively, tangent-view approaches [18, 19, 20] project panoramas onto multiple views, estimate per-view depth, and spatially composite patches into panoramic depth maps. Vertically compressed methods [24, 25, 26] adapt layout estimation techniques, reducing panoramas to 1-pixel height and employing Bi-LSTMs or self-attention for depth estimation.
II-B Bi-Projection Inputs
Methods using bi-projection inputs project panorama images onto two distinct perspectives. Of the bi-projection input methods, equirectangular projection is usually used as the primary perspective by default. These methods usually use shared or dedicated branch networks to predict corresponding depth maps. Then, the another perspective-specific depth estimate are reprojected to the equirectangular domain and composited with directly predicted equirectangular depth maps to yield refined depth estimations. Representative examples include BiFuse[12], BiFuse++[27], and UniFuse[28], which project onto cube maps; HRDFuse[29] and GA360Fuse[30], which fuse equirectangular and tangent depth predictions; and Elite360D[31], which fuses equirectangular and ICOSAP[22] perspectives for enhanced accuracy. Intuitively, such methods seem better suited to handle distortion than single-view panoramic depth estimation. However, current public dataset rankings show that supervised learning on pure equirectangular images achieves superior performance. Moreover, dual-view projection inevitably introduces significant additional computational overhead during training and inference. Therefore, the benefits of this approach may not justify its computational cost.
II-C Background Based Methods
Current state-of-the-art methods demonstrate strong quantitative performance across benchmarks. However, their 3D depth visualizations frequently exhibit structural inaccuracies in room geometry and over-smoothed corners. To address this limitation, recent approaches[32] integrate room structural priors into panoramic depth estimation. The core challenge lies in efficiently computing accurate background depth maps, which require precise room layout estimation.
Indoor panoramic layout estimation aims to detect room wall boundaries from input panoramas. Existing research predominantly adopts the Manhattan World assumption. LayoutNet[33] directly predicts per-pixel corner and boundary probability maps. Conversely, Dula-Net[34] decodes panoramas into equirectangular and perspective views, extracting wall-floor/ceiling boundaries via semantic segmentation. EquiConv[35] employs specialized convolutions to generate corner/edge probability maps. These early methods share a 2D segmentation paradigm, outputting full-resolution probability maps—an inefficient approach given layout estimation only requires boundary/corner identification. HorizonNet[36] thus reformulates the task as 1D sequence prediction using Bi-LSTMs, while HoHoNet[29] employs multi-head self-attention for similar 1D representation. LED-Net[37] further predicts layouts from horizontally compressed representations. Subsequently, AtlantaNet[38] projects panoramas onto ceiling/floor planes to avoid occlusion, predicting layouts via amalgamated projections. DMH-Net[39] extends this by mapping to cubemap faces before predicting boundaries and corners.
A critical secondary challenge involves fusing background depth with panoramic depth predictions. The recent BGDNet[32] incorporates room geometry as background depth when predicting the final depth map. However, it needs pretrained HorizonNet and SAM[40] to extract the background depth map from input panorama image. In this paper, we adopt a multi-task learning approach to integrate room layout and segmentation predictions needed for background depth calculation within a single framework. We also design a background depth calculation strategy based on background segmentation and a fusion strategy for coarse and background depth guided by background segmentation weights. Our method is detailed in the following sections.
Figure 2: The structure diagram of our proposed room geometry guided depth estimation framework. In terms of model structure, our framework includes a shared panorama encoder and three task-corresponding decoders. Based on the obtained layout map, coarse depth map, and background-segmentation map, the proposed framework decode fine-grained depth prediction.
III Methodology
The room structural regularized depth estimation framework proposed in this paper is shown in Fig.2. From the view of neural network architecture, the framework includes a shared panorama encoder and three task-corresponding decoder modules. Moreover, our framework contain a room geometry-based background depth resolving strategy and a background-segmentation-guided fusion mechanism. The panorama encoder extract the multi-stage feature sets $F=\{f^{2^{i+1}\times}_{i}|i=1,2,3,4\}$ from the panorama image. The decoder modules corresponding to the subsequent three tasks use these features as input to estimate the corresponding results. The proposed room geometry-based background depth resolving strategy calculate the background depth based on panorama depth, room layout, background segmentation predictions. Then, background-segmentation-guided fusion mechanism regularize the depth decoder’s prediction based on background depth map and background segmentation predictions. In the following subsections we will introduce each module in detail.
III-A Architecture
Figure 3: The structure diagram of layout feature aggregation module and pixel feature aggregation module.
Panorama Encoder: In this paper, the framework we proposed uses the backbone proposed by PanoFormer as a common feature encoder. The backbone consists of the panorama transformer block and convolution 2D layer based downsample layer. The panorama transformer block contains a window self-attention mechanism designed for the panorama imaging process and a feed-forward layer designed based on depthwise separable convolution[41]. The entire backbone consists of an input projection layer and $4$ stage blocks. The input projection layer consists of a convolution layer whose kernel size is $3\times 3$ with stride of $2$, a 2D batchnorm layer and a ReLU activation function. The subsequent $4$ stage blocks all consist of a panorama transformer block and a double-downsampling layer composed of a 2D convolutional network. Each stage block will output the corresponding double-downsampled feature map result.
Layout Decoder: Considering that layout estimation is a non-pixel-level task while depth estimation and background segmentation are both pixel-level tasks, and some existing work [42] also mentioned that different tasks may have different requirements for features of different scales. To this end, we design a specific layout FA module for the layout task to further process the features extracted by the panorama encoder. The structure diagram of the layout feature aggregation module for the layout task is shown in Fig.3 (a) . This module use the height compress convolution with spatial stride $(2,1)$ to compress the feature maps $f^{4\times}_{1},f^{8\times}_{2},f^{16\times}_{3},f^{32\times}_{4}$. Then the compressed feature map is flattened to obtain a feature vectors set $\hat{f}_{1},\hat{f}_{2},\hat{f}_{3},\hat{f}_{4}$. We then cat these flattened features together and, following the Hohonet[25] model design experience, use a multi-head self-attention module to extract global feature information from the flattened feature vectors and output the final feature vector $\hat{f}_{cat}$. Based on the output features of the layout FA module, the layout decoder uses boundary and corner heads, consisting of a convolution 1D layer, a ReLU layer, and a batchnorm 1D layer, to predict the final room layout $S_{room}$.
Depth Decoder: Considering that both depth estimation and background segmentation tasks are pixel-level tasks, we follow the design idea of Hohonet and let them use a common pixel-level feature aggregation to process the multi-stage panorama feature. The structure of the pixel-level feature aggregation module we use is shown in Fig. 3 (b). This module consists of a panorama transformer block and an upsampling module. Given multi-scale feature map set $f^{1}_{4\times},f^{2}_{8\times},f^{3}_{16\times},f^{4}_{32\times}$, this module’s feature integration process can be described as:
$\hat{f}=\prod^{3}_{i=1}[UP(PT(f^{2^{i}\times}_{i}))+f^{2^{i+1}\times}_{i+1}]$ (1)
On the integrated multi-scale feature $\hat{f}$, the depth estimation decoder predict the coarse depth maps $S^{p}_{depth}=\{d_{ij}|i=1,2,...,W;j=1,2,...,H\}$.
background-Segmentation Decoder: As mentioned above, the background segmentation task and the depth estimation task share a pixel-level feature aggregation module. The structure diagram of the pixel-level feature aggregation module is shown in Fig. 3(b). Therefore, in our framework, the background-segmentation decoder itself has only one prediction head consisting of a convolution 2D layer, a ReLU function, and a batchnorm 2D layer. This prediction head predicts the corresponding 2D background segmentation result $S_{seg}=\{0\leq p_{ij}\leq 1|i=1,2,...,W;j=1,2,...,H\}$ based on the integrated multi-scale feature $\hat{f}$.
III-B Room Geometry-based Background Depth Resolving Strategy
Figure 4: P is the camera center, A and B are the upper and lower boundary points of the wall corresponding to point P, and D is an arbitrary point on the wall plane. The lengths of AB in the image are known, and the corresponding angles $\phi_{c}$ and $\phi_{f}$ can be calculated based on spherical camera geometry. Based on the depth of point P predicted by the depth decoder, $d_{c}$ and $d_{f}$ can be calculated.
Given a room layout estimation prediction $S_{room}\in R^{W\times 3}$ to resolve the corresponding background depth, we need to obtain the camera-to-ground distance information. Existing methods often assume a fixed camera height when calculating background depth based on layout information. This assumption is stable and reliable in virtual rendering datasets, but it is not necessarily reliable when collecting data in the real world using a panoramic depth camera. To obtain more stable and reliable camera height information for each scene, we first designed a camera height resolution strategy.
As shown in the Fig. 4, define the coordinates of the camera origin $O$, and draw a normal line from the camera plane to the image plane, which intersects the wall plane at point $P$. Then draw a straight line perpendicular to the camera plane at point $P$, which intersects with the ceiling-wall boundary line in the room layout prediction at point $A$, and intersects with the floor-wall boundary line at point $B$. What our camera height calculation strategy needs to do is to calculate the camera height in the Fig. 4. Our method use the predictions of the depth decoder and room layout decoder as the input of camera height calculating method. The room layout predictions are responsible for providing the image plane coordinates of the intersection points $A$ and $B$ in the panorama image and depth maps. Based on the results of the depth maps prediction, we can get the distance between the corresponding line segments $OA$ and $OB$. On this basis, the distance from the camera to the ceiling and the camera height can be conveniently calculated using the following formula:
$|AP|=d_{c}\times sin(\phi_{c})$ (2)
$|PB|=d_{f}\times sin{\phi_{f}}$
The corresponding angles $\phi_{c}$ and $\phi_{f}$ can be calculated using the geometric imaging principle of a panoramic camera. Assuming that the vertical coordinate of point $A$ is $u_{ceil}$ and the vertical coordinate of point $B$ is $u_{floor}$, the calculation formulas for the two angles can be obtained using the following formulas:
[…]
After determining the camera height and the distance between the camera and the ceiling, we can use the wall information from the room layout predictions to calculate the corresponding background depth maps for each pixel belonging to the background wall. The depth maps for the ceiling and floor can be simply calculated based on the angles calculated from the pixels using the following formula:
$d^{i}_{floor}=\frac{|PB|\times H}{(u^{i}_{floor}-0.5H)\times\pi}$ (4)
$d^{i}_{ceil}=\frac{|AP|\times H}{(0.5H-u^{i}_{ceil})\times\pi}$
For any point $D$ on the wall plane that is tilted toward the floor, assume its equirectangular plane coordinates on the panorama image are $(u_{i},v_{i})$. First, we can calculate the horizontal and vertical offset angles $\rho_{D}$ and $\phi_{D}$ of this point relative to the camera center $P$ using the following formulas:
[…]
$\rho_{D}=(1-\frac{v_{i}}{W})\times\pi$
Based on the two deflection angles, to obtain the depth of $D$, $d_{D}=|OD|$, we first draw a perpendicular line through point $D$ to line segment $AB$ and intersect $AB$ at point $E$. First, based on the vertical offset angle $\phi_{D}$, we can obtain $|OE|=cos(\phi_{D})|OP|$, where the value of $|OP|$ can be directly obtained from depth prediction. Then, based on the horizontal offset angle $\rho_{D}$ and $|OE|$, we can simply solve $d_{D}=cos(\rho_{D})|OE|$. Similarly, any point on the wall plane that is deflected toward the ceiling can also use a similar method to solve for its corresponding depth value. Finally, combining all the solved depth values, we can obtain $S_{back}$.
III-C Background-segmentation-guided Fusion Mechanism
As mentioned above, based on the corase depth map and the background depth map, our framework needs to fuse the predicted depth $S^{p}_{depth}$ and the background depth $S_{back}$ to obtain an accurate final depth estimation result $S^{final}_{depth}$. Considering that $S_{back}$ is naturally suitable as an upper bound for depth estimation when the layout prediction is relatively accurate. Then $S_{back}$ as an upper bound can be used to ensure that the depth of pixels originally belonging to the rescue, ceiling, ground and other areas do not exceed the area of the room itself. The depth value of the pixel belonging to the foreground object area can be as close to $S^{p}_{depth}$ as possible. Therefore, in the process of fusing the two depth maps, we need a weight to determine whether the current pixel is an object belonging to the foreground area or the background area. The probability map $S_{seg}$ predicted by the background-segmentation task in our framework can just be used as such a fusion weight. Therefore, the entire fusion process can be described as the following formula:
$S^{final}_{depth}$ $=\{d^{final}_{ij}|i=1,2,.,W;j=1,2,.,H\};$ (6)
$etl.$ $\ d^{final}_{ij}=d^{back}_{ij}\times p_{ij}+d^{p}_{ij}\times(1-p_{ij});$
$p_{ij}\in S_{seg};$
$d^{back}_{ij}\in S_{back};$
$d^{p}_{ij}\in S^{p}_{depth};$
III-D Objective Function
[…]
When training the layout branch, we refer to the settings of HorizonNet and use the binary cross entropy and L1 loss function to calculate. When training the depth estimation branch, we used the same Huber[43] (or Berhu[44]) loss and gradient loss[45] as the baseline Panoformer. The hyper-parameters within the objective functions of these two branches were also consistent with those of HorizonNet and PanoFormer.
Considering that existing panoramic image datasets usually do not contain background segmentation annotations. When training the background segmentation decoder, we first use the background depth map $S_{back}$ predicted by the framework and the corresponding ground-truth depth map $S^{gt}_{depth}$ to calculate the corresponding background segmentation label $S^{gt}_{seg}$. The specific solution formula is as follows:
$D^{res}=|S^{gt}_{depth}-S_{back}|;$ (7)
$S^{gt}_{seg}=\{p^{gt}_{ij}=\Gamma(d^{res}_{ij}<\gamma)|d^{res}_{ij}\in D^{res}\}$
where $\Gamma$ is a binary function, which takes the value 1 when the condition in the brackets is true and takes the value 0 when it is false.
When optimizing the partial network for the background-segmentation task, we use a simple focal loss as our object function, which is calculated as follows:
$L_{seg}=-\frac{1}{M}\sum^{M}_{u=1}\frac{1}{H*W}\sum^{W}_{i=0}\sum^{H}_{j=0}\alpha(1-\hat{p}_{ij})^{\eta}log(\hat{p}_{ij});$ (8)
where $\hat{p}_{ij}=p_{ij}$ if $p^{gt}_{ij}=1$ else $\hat{p}_{ij}=1-p_{ij}$. Following the common setting of focal loss, we set the $\alpha$, $\beta$ as 0.5, 2.0 relatively.
In summary, the overall object function of our proposed framework can be described as:
$L_{all}=\lambda_{1}*L_{layout}+\lambda_{2}*L_{depth}+\lambda_{3}*L_{seg}$ (9)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are three manual setting parameters.
IV Experiment
IV-A Experiment Setting
In this section, we introduce the experiments to verify the proposed depth estimation framework. In our experiments, we selected three datasets: Stanford2D-3D[46], MatterPort3D[47], and Structure3D[48]. Stanford2D-3D and MatterPort3D are two datasets collected in the real world, while the Structure3D dataset is a synthesized dataset. Real-world datasets contain a large amount of noise points introduced during the acquisition process as described in OnimiDepth[13]. This noise introduces local jumps that severely hinder model performance. Therefore, before training models on Stanford2D3D and Matterport3D, we used a proposed layout-based dataset denoise strategy to constrain the depth of these noise points to within the room.
Stanford2D-3D: The Stanford2D-3D dataset contains 1413 panoramic images collected from 3 types of buildings divided into 6 large areas. We follow the official practice to divide the dataset into training and test sets, and downsample all depth maps and RGB images to $512\times 1024$ size images.
Matterport3D: Matterport3D contains 10,800 panoramic images collected from 90 different rooms. The camera used in the collection process of this dataset is Matterport’s Pro 3D camera. In this part of the dataset, we also use 61 room images for training and 29 room images for testing. All RGB and depth images are also downsampled to $512\times 1024$ size.
Structured3D: The Structured3D dataset contains 196K rendered panoramic images and corresponding depth maps, covering 12,835 rooms in 3,500 scenes. Each room is created manually using CAD models of furniture, which are in real-world dimensions and used in real production. In this dataset, we follow the official recommended setting [48], using the data of the first 3000 scenes as the training set, the data of 3000-3250 scenes as the validation set, and the data of the last 250 scenes as the test set.
Metric: Following the previous works, we use some standard evaluation metrics, which include: relative error (abs rel), squared relative error (sq rel), root mean squared error (rmse), and three threshold percentage $\delta<\varsigma^{t}(\varsigma=1.25,t=1,2,3)$ denoted as $\delta_{t}$.
Dataset denoise strategy: Considering that datasets collected in the real world contain a large amount of noisy points, we have also designed a denoising strategy for panoramic depth map datasets. As mentioned earlier, our method uses the outputs of the layout decoder and depth decoder to calculate the corresponding background depth map. Here, we use this calculated background depth map to constrain the dataset collected in the real world. First, for areas in the real-world depth ground-truth maps where measurement fails, we directly replace them with the background depth value. Then, our method sets a threshold to judge the ground-truth depth. If the depth point in the ground-truth dataset of the original data is converted to 3D space and its coordinates are outside the room and the distance from the room wall is more than 1 meter, we identify the pixel as a noise pixel. For these pixels, we also use the depth in the background depth map as a replacement.
Training setting: We use Adam as the optimizer, and the parameters of the optimizer are basically the basic settings of the pytorch framework. For the learning rate scheduling strategy, we choose one-cycle[49], set the initial learning rate to 0.0001, and the minimum learning rate to 0.0000001. Our hardware experimental platform is configured with AMD Epyc 7003 CPU and 4-card RTX 4090 GPU. During the training process, we set the batch size on each card to 2. In the data set enhancement part, we used random horizontal angle rotation and random horizontal flipping with reference to Panoformer[10]. For horizontal angle rotation, we set the interval of random angle to $\frac{\pi}{4},\frac{\pi}{4}]$.
Parameter setting: The training objective function used by our framework is shown in Section 4.4, which contains three subcomponents: $L_{front}$, $L_{depth}$, and $L_{layout}$. Considering that some datasets do not contain complete layout annotations, and our framework is extremely dependent on the accuracy of layout estimation, we will first extract all layout annotations from the three selected datasets to form a large layout dataset to pretrain the layout task. Then, when training the depth estimation task on specific datasets, we decide whether to finetune the weights of the layout and feature encoder parts based on whether the dataset contains complete layout annotations. Specifically, the Stanford2D-3D data does not provide complete layout annotations, so we set the corresponding hyper-parameter $\lambda_{1}$ to 0 when training our framework, and the corresponding $\lambda_{2},\lambda_{3}$ to 1.0 and 0.4. The Matterport3D and Structure3D datasets provide complete layout annotations, so we set $\lambda_{1},\lambda_{2},\lambda_{3}$ to 0.01, 1.0, and 0.4, respectively. In addition, some hand-tuned hyper-parameter settings related to this study, such as $\rho_{1}$ and $\rho_{2}$ in equation was set to 1.0 and 0.5. Other hyper-parameters that are not relevant to this study are set according to the common specifications in the current field.
IV-B Performance Comparison
Figure 5: The 3D visualize results of final depth estimation. For each scene, we selected three perspectives: top view, side view, and internal perspective to display the three-dimensional visualization effect of the point cloud converted from the depth map.
IV-B1 Quantification Comparison with SoTA Depth Estimation Methods
TABLE I: Quantification comparison with State of the Art depth depth estimation methods on Matterport3D
Dataset Method Pub’ Year Classic Metrics
$\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE MRE MAE
Matterport3D EGFormer[11] ICCV 2023 0.8158 0.9390 0.9735 0.6025 0.1517 0.1473
OminiFusion[19] CVPR 2022 0.9040 0.9757 0.9919 0.4261 0.0552 0.0900
Bifuse[12] CVPR 2020 0.8452 0.9319 0.9632 0.6295 0.2408 0.3470
UniFuse[28] IEEE RAL 2021 0.8897 0.9623 0.9831 0.4941 - 0.2814
HRDFuse[29] CVPR 2023 0.9162 0.9669 0.9844 0.4433 0.0936 0.0967
RSDNet[50] CVPR 2022 0.443 0.097 0.248 0.906 0.971 -
SN360[51] IEEE Access 2025 0.4483 - - 0.9392 0.9808 0.9932
GLPanoDepth[23] IEEE TIP 2024 0.8641 0.9561 0.9808 0.5223 - 0.2998
Panoformer[10] ECCV 2022 0.9184 0.9804 0.9916 0.3635 0.0571 0.1013
SGFormer[15] IEEE TCSVT 2025 0.8946 0.9642 0.9859 0.4790 - 0.2748
Ours 0.9199 0.9820 0.9984 0.2436 0.0425 0.0829
TABLE II: Quantification comparison with State of The Art depth estimation methods on Stanfor2d3d
Dataset Method Pub’ Year Classic Metrics
$\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE MRE MAE
Stanford2d3d EGFormer[11] ICCV 2023 0.8185 0.9338 0.9736 0.4974 0.1408 0.1528
OminiFusion[19] CVPR 2022 0.8940 0.9714 0.9900 0.3715 0.0543 0.0961
Bifuse[12] CVPR 2020 0.8660 0.9580 0.9860 0.4142 0.1209 0.2343
UniFuse[28] IEEE RAL 2021 0.8711 0.9664 0.9882 0.3691 - 0.2082
HRDFuse[29] CVPR 2023 0.8941 0.9778 0.9923 0.3452 0.0503 0.0984
FreDSNet[52] 2022 0.8424 0.9583 0.9863 0.2727 0.0952 0.1327
RSDNet[50] CVPR 2022 0.394 0.098 0.209 0.903 0.974 -
SN360[51] IEEE Access 2025 0.2917 - - 0.9369 0.9846 0.9942
GLPanoDepth[23] IEEE TIP 2024 0.9015 0.9793 0.9901 0.3493 - 0.1932
Panoformer[10] ECCV 2022 0.9394 0.9838 0.9941 0.3083 0.0405 0.0619
SGFormer[15] IEEE TCSVT 2025 0.8998 0.9693 0.9908 0.3406 - 0.2017
Ours 0.9479 0.9857 0.9943 0.2359 0.0285 0.058
TABLE III: Quantification comparison with State of the Art depth estimation methods on Structured3D
Dataset Method Pub’ Year Classic Metrics
$\delta_{1}$ $\delta_{2}$ $\delta_{3}$ RMSE MRE MAE
Structure3d EGFormer[11] ICCV 2023 0.7979 0.9071 0.9455 0.6841 0.4509 0.2205
OminiFusion[19] CVPR 2022 0.6921 0.8831 0.9501 0.4951 - 0.2981
Bifuse[12] CVPR 2020 0.8594 0.9400 0.9672 0.5213 0.2455 0.1573
UniFuse[28] IEEE RAL 2021 0.8542 0.9399 0.9676 0.5016 0.2319 0.1506
HRDFuse[29] CVPR 2023 0.7561 0.9161 0.9631 0.4061 - 0.2451
BGDNet[32] CVPR 2024 0.8336 0.9377 0.9731 0.3490 - 0.1656
SGFormer[15] IEEE TCSVT 2025 0.9613 0.9896 0.9957 0.2429 - -
PanoFormer[10] ECCV 2022 0.8943 0.9536 0.97431 0.3017 0.1201 0.1546
Ours 0.9679 0.9907 0.9983 0.1935 0.0414 0.0613
Figure 6: Visualization of background depth map guided dataset denoising effect diagram.
[…]
The Stanford2d3d dataset is a relatively early dataset, and various existing depth estimation methods have achieved relatively good performance on this dataset. However, on this dataset, the noise processing strategy and multi-task depth estimation framework proposed in this paper still show excellent effectiveness. Specifically, our method significantly outperforms the current SOTA method by 3.68 percentage points in the RMSE, the evaluation indicator that the industry attaches the most importance to.
At the same time, in MRE and MAE, which are both numerical indicators, the performance advantage of our method over the existing methods is not as obvious as RMSE, but it is still intuitive, which further proves the robustness and superiority of the method proposed in this paper. In the three indicators $\delta_{1},\delta_{2}$ and $\delta_{3}$ at the ratio level, our method also has a leading advantage, but this advantage is not as obvious as the numerical indicators. We believe that this is because the existing methods have reached a very high level in the three indicators at the ratio level, so it is more difficult to achieve higher indicators.
On the Matterport3D dataset, our method also achieved the current best performance in RMSE, the core evaluation indicator of depth estimation, which fully demonstrated the effectiveness of the proposed method and data processing strategy in key accuracy metrics. However, on the relative error threshold indicator ($\delta<$ 1.25), our method failed to achieve the best results. We believe this is related to the introduced layout constraints: this constraint tends to optimize the absolute error between the prediction and the true value, but in areas with small true depth values, this optimization may result in a limited range of change in the predicted value, making it more difficult to meet the strict 1.25 times relative error requirement. Despite this, we still achieved a level close to SOTA on this indicator. Importantly, this method is significantly ahead of existing work in both the core rmse and rmslog indicators.
On the Structured3d dataset, we mainly compare with BGDNet, which uses a similar idea. The results released in the BGDNet paper are trained on the replica and then val on the structure3d dataset, while our method is trained on Structured3d and then validated on the corresponding validation set. This comparison may not be fair, but considering that the BGDNet paper is not open source, we can only use this method for comparison. Due to the differences in the settings of train and val, the method proposed in this paper has a very obvious advantage in the six indicators of concern.
TABLE IV: Quatitative analysis of our proposed dataset denoise method.
Dataset Setting Methods RMSE MRE MAE
Without Denoise PanoFormer 0.3083 0.0405 0.1013
Our Methods 0.2673 0.0375 0.0912
Denoised Dataset PanoFormer 0.2872 0.0392 0.0892
Our Methods 0.2359 0.0285 0.0580
TABLE V: The training impact of each task decoder to depth decoder
Depth
Decoder
Front-Seg
Decoder
Layout
Decoder
RMSE MRE MAE
0.3635 0.0571 0.1013
0.3621 0.0531 0.1011
0.3656 0.0582 0.1101
0.3592 0.0511 0.1002
TABLE VI: The accuracy of resolved background depth maps.
Methods RMSE MRE MAE
PanoFormer’s Background Region 0.3214 0.0623 0.1123
Decode from layout prediction 0.0043 0.0032 0.0023
Decode from layout ground-truth 0.0001 0.0001 0.0001
TABLE VII: The comparison of our proposed fusion methods and BGDNet
Methods RMSE MRE MAE
Baseline 0.3635 0.0571 0.1013
Fusion BGDNet 0.2823 0.0498 0.0922
Fusion ours 0.2436 0.0425 0.0829
Figure 7: The experiment results of explore the inference of hyper-parameter $\lambda_{2}$ and $\lambda_{3}$. The x-axis in this figure represents the value of the corresponding hyper-parameter setting, while y-axis represents the RMSE value of our proposed framework’s depth prediction.
IV-B2 Visualize Comparison with SoTA Depth Estimation Method
In this section, we perform visualization operations on the scene as shown in the figure. In this set of experiments, we selected the baseline panoformer in this paper as the comparison object. It is worth noting that in this set of experiments, panoformer and the depth estimation framework proposed in this paper both use denoised datasets during training. In this set of experiments, we selected several representative scenes on the validation set for demonstration. From the corresponding 3D visualization results, it can be seen that the depth estimated by our method is significantly more accurate in corners and walls. At the same time, thanks to the prior knowledge of room layout introduced by us, the 3D visualization of the prediction results of our method is more reasonable in terms of overall structure. At the same time, due to the existence of our dataset denoising strategy, panoformer and our method do not have the phenomenon of overfitting of noise points in patches as shown in Fig. 1. The existence of these phenomena fully demonstrates the effectiveness of the depth estimation method we proposed.
IV-C Visualize and Quatitative analysis of Dataset Denoise Method
IV-C1 Visualize Analysis of Denoise Method
We conducted a visualization experiment as shown above for the panoramic depth images of the 2D-3D-S dataset collected in the real world. After mapping each pixel in the panoramic depth images to 3D space, we found that there are a large number of noise points in each scene. For these noise points, the processing strategy of the existing panoramic depth estimation method is to directly preset a maximum distance threshold. In the 2D-3D-S dataset, this threshold is set to 10 meters. The depth values of all areas greater than this threshold will be artificially set to 0. On the one hand, this processing method destroys the correlation between the pixels of the original panoramic depth map, thereby affecting the learning process of the model. On the other hand, it will also cause some areas in the scene with depth values less than 10 but should fall inside the room to fall outside the room. In order to preserve the pixel correlation of the processed panoramic depth map and prevent the area falling inside the room from falling outside the room, we proposed a strategy to preprocess the indoor panoramic depth map using pretrained layout. In this strategy, we first collected as many parts of the existing dataset that contain layout annotations as possible to form a layout dataset of a relatively reasonable size to pretrain the layout estimation part of the framework proposed in this paper. Considering that the layout prediction branch can already obtain a high 3D IoU in our experiments, we directly use this part of the prediction results to process the depth estimation dataset in the actual processing process. In our proposed processing strategy, we first use the layout prediction branch to estimate the layout of the current room, and then use the panorama depth ground-truth to count the height of each wall of the room, and use this height result to solve the 2D layout into the corresponding background depth map. Finally, we use this background depth map to constrain the original panorama depth ground-truth as shown in Fig. 6. In this constrained process, we use the background ground-truth as a mask to ensure that the depth of the background region does not exceed the range of the room itself.
IV-C2 Quatitative Analysis of Denoise Method
We use the Stanfor2d3d dataset as the object to verify the effectiveness of the proposed noisy image processing strategy. In the experiment, we use our proposed framework and PanoFormer to perform training and val on the processed datasets before and after processing. The experimental results are shown in Table IV. From the results in Table IV, the performance of PanoFormer and our proposed framework trained on the processed dataset is significantly higher than that trained on the original dataset, which fully demonstrates the performance improvement brought by our data processing strategy.
IV-D Ablation Study
IV-D1 The Effect of Framework’s Each Component
Figure 8: The visualize of Background segmentation. The left column of the figure is input panorama RGB images, the middle column is background segmentation predictions, the right column is corresponding ground-truth.
Figure 9: The visualize of room layout estimation. The first column shows the input RGB image. The second column shows the room layout estimation results displayed on a 2D equirectangular image. The top row in each 2D equirectangular sub-image represents the probability of a corner point in that column. The third and fourth columns show the corresponding 3D bird’s-eye views of the prediction and layout ground-truth, respectively.
In this section, we analyze the impact of each proposed component on the whole. Considering that we proposed a framework for auxiliary-task assisted depth estimation tasks, we mainly discuss two aspects of the impact of each component: 1) The impact of auxiliary-task decoder for depth estimation decoder; 2) The effect of our proposed background depth resolving and fusion strategy. The experimental dataset we chose is Matterport3D because it contains complete layout estimation annotations.
Regarding the aspect 1), the corresponding experimental results have been put into Table V. It is worth noting that the loss weight we used in this group of experiments is the combination verified in subsequent section. From the results, it can be seen that the auxiliary-task decoder has very little impact on the depth decoder during training. This is primarily because we intentionally set a low loss weight for the auxiliary-task decoder. This design stems from two considerations: first, optimizing each task in a multi-task learning framework requires extensive experimentation and manual tuning. Second, and more importantly, our framework’s primary objective is to use the auxiliary task’s output to constrain the depth decoder directly, rather than to improve it indirectly by learning more generalized features in the shared encoder.
Regarding the aspect 2) we want to discuss, we conducted corresponding experiments and put them in Table VI and Table VII. The experimental results of the background depth resolving strategy are shown in Table VI, and the experimental results of the fusion strategy used are shown in Table VII. When evaluating the background depth resolving strategy, we used the depth ground-truth provided by the Matterport3D dataset as a benchmark. When evaluating the resolved background depth, we used the ground-truth of the background segmentation as a mask to ensure that only the background region is evaluated during the eval process. Judging from the results in Table VI, our background depth resolving strategy is more accurate than the deep learning model. In terms of fusion strategy, we chose BGDNet[32] as the comparison object. In this set of experiments, the configuration of the fusion strategy used by BGDNet is consistent with the original paper. The experimental results demonstrate that incorporating background depth as a constraint effectively enhances depth estimation performance. Moreover, our proposed fusion strategy yields even more substantial improvements compared to that of BGDNet.
In summary, our discussion validates both the proposed framework and the collaborative design of its subtasks.
IV-D2 Analysis of loss weights
The impact of different loss weights hyper-parameters of the proposed framework has been shown in Fig. 7. Here, we use the MatterPort3D to conduct experiments. As the layout decoder in our framework has been pretrained, we simply fix its $\lambda_{1}$ to 0.01. For $\lambda_{2}$ and $\lambda_{3}$, we first fix the $\lambda_{3}$ to 1.0 and search the best setting of $\lambda_{2}$. Then, fix the $\lambda_{2}$ to the searched results and search the setting of $\lambda_{3}$. The Fig. 6. show that on the MatterPort3D dataset, when we set $\lambda_{2}=1.0$ and $\lambda_{3}=0.4$, the proposed framework can achieve the best performance.
IV-E Visualize Analysis of Background Segmentation and Layout Estimation
In this section, we visualize the prediction results of the front-segmentation and layout estimation task. We used the test set data of Structured3D to perform correspond experiment. The specific visualization results are shown in Fig. 8 and Fig. 9.
Fig. 8 shows the visualization results of the background segmentation task. In our framework, the prediction results of the background segmentation task are used as weight to fuse the coarse prediction and background depth. When visualizing the results of background segmentation, we mapping the prediction score maps from $[0,1]$ to $[0,255]$ and save them as gray images. Judging from the score maps shown in Fig. 8, our background branch can accurately distinguish between the background and background areas in the input image. However, in some areas that are very close to the wall and have a similar depth to the background area, the discrimination ability of our background branch still needs to be improved.
Fig. 9 shows the estimation results of our layout estimation branch. From the 2D visualization of the prediction results, we can see that the edge areas and corner points of the wall can be identified relatively accurately. However, in some areas far away from the camera, the prediction accuracy of the corner points still needs to be improved. In addition to these issues, another point worth noting is that the current room layout estimator assumes that the walls of the room are all cubic structures. This makes our method perform worse than the normal depth estimator when estimating the depth of some rooms with curved wall structures. We consider conducting more detailed research on this issue in future work.
V Conclusions
This paper has proposed a panoramic depth estimation framework with room geometry constraints. The framework employs multi-task learning, where a shared encoder extracts features that are decoded into three outputs: coarse depth maps, background segmentation masks, and room layouts. These predictions enable our method to initially reconstruct background depth from the layout information. Subsequently, background depth and coarse depth are fused using the background segmentation mask as a weighting mechanism, ultimately generating the final depth prediction. Extensive experiments on real-world and synthetic datasets demonstrate significant performance improvements over current methods.